OpenStudy (anonymous):

x/5 + x/3 =24 What is X? ~is it 5~

jimthompson5910 (jim_thompson5910):

first combine the fractions on the left side x/5 + x/3 =24 3x/15 + x/3 =24 3x/15 + 5x/15 =24 (3x + 5x)/15 = 24 8x/15 = 24 do you see what to do from here?

OpenStudy (anonymous):

so its not 5...

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

8x/15 = 24 8x = 24*15 8x = 360 x = ???

OpenStudy (anonymous):

45

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you, i understand it better now :)

jimthompson5910 (jim_thompson5910):

I'm glad you do